The length of a rectangular garden is 8 feet longer than its width. If the garden's perimeter is 184 feet, what is the area of the garden in square feet?
Answer by rfer(16322)

let l=w+8=50
let w=w=42
2(w+8)+2w=184
2w+16+2w=184
4w=168
w=42
----------------
a=lw
a=50*42
a=2100 sq ft.